OpenStudy (anonymous):

does not understand how -5/12 - 5/6= 5/4

jimthompson5910 (jim_thompson5910):

\[\Large -\frac{5}{12}-\frac{5}{6}\] \[\Large -\frac{5}{12}-\frac{5}{6}\times \frac{2}{2}\] \[\Large -\frac{5}{12}-\frac{5 \times2}{6 \times 2}\] \[\Large -\frac{5}{12}-\frac{10}{12}\] \[\Large \frac{-5-10}{12}\] \[\Large -\frac{15}{12}\] \[\Large -\frac{5}{4}\] So \[\Large -\frac{5}{12}-\frac{5}{6} = -\frac{5}{4}\]

OpenStudy (anonymous):

in fractions addition or subtraction , we have to get a common denominator , and in your question : -5/12 - 5/6 , we have different denominators right?